OpenStudy (anonymous):

Which of the following is equivalent to log base 9 27 ? Which of the following is equivalent to log base 9 27 ? @Mathematics

OpenStudy (asnaseer):

remember that if:\[\log_a{x}=y\]then this implies:\[a^y=x\]

OpenStudy (asnaseer):

you need to state what options you have been given

OpenStudy (anonymous):

1.452.....2/3...3....3/2

OpenStudy (asnaseer):

lets say:\[\log_9{27}=x\]then:\[9^x=27\]so:\[(3^2)^x=27\]\[3^{2x}=27=3^3\]so:\[2x=3\]\[x=\frac{3}{2}\]